what is the value of
-3+4a=-11
2x+6(3x-7)=-2
4-3(x+1)=2x+1

Answer by ankor@dixie-net.com(22740)

what is the value of
-3 + 4a = -11
add 3 to both sides
4a = -11 + 3
4a = -8
a = -8/4
a = -2
Check this for yourself in the original equation
:
2x + 6(3x-7) = -2
distribute the 6
2x+ 18x - 42 = -2
20x = -2 + 42
20x = 40
x = 40/20
x = 2
Do the same for this one
:
4 - 3(x+1) = 2x + 1
4 - 3x - 3 = 2x + 1
-3x + 1 = 2x + 1
-3x - 2x = 1 - 1
-5x = 0
x = 0
Check this in the original equation
4 - 3(0+1) = 2(0) + 1
4 - 3 = 0 + 1